a particle is projected with a speed of 24m/s.find the two angles of projection to give a range of 14m
 

Grade:10

1 Answers

Vikas TU
6 years ago
Dear Student,
For angular projection, we have R = (u2sin2θ)/g.
To get R = 14 m with θ = 24 m/s, we have 14 = 576 sin2θ/9.8
sin2θ = 14 x 9.8/576.
2θ = sin-1 (14 x 9.8/576) = 13.78 degrees.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)